# Insect-Scale Tailless Robot with Flapping Wings: A Simple Structure and Drive for Yaw Control

### v1

@@ Abstract
Insect-scale micro-aerial vehicles, especially, lightweight, flapping-wing robots, are becoming increasingly important for safe motion sensing in spatially constrained environments such as living spaces. However, yaw control using flapping wings is fundamentally more difficult than using rotating wings. In this study, an insect-scale, tailless robot with four paired tilted flapping wings (weighing 1.52 g) to enable yaw control was fabricated. It benefits from the simplicity of a directly driven wing actuator with no transmission and a lift control signal; however, it still has an offset in the lift force. Therefore, an adaptive controller was designed to alleviate the offset. Numerical experiments confirm that the proposed controller outperforms the linear quadratic integral controller. Finally, in a tethered and controlled demonstration flight, the yaw drift was suppressed by the wing-tilting arrangement and the proposed controller. The simple structure drive system demonstrates the potential for future controlled flights of battery-powered, tailless, flapping-wing robots weighing less than 10 grams.
I INTRODUCTION
Unmanned aerial vehicles (UAVs) are beginning to be used for extensive monitoring and inspection of farms, roads, buildings, and bridges. However, because they generate their lift using rotating propellers, their movement in spatially constrained environments, such as living spaces, poses safety risks in the event of collisions. Therefore, micro-aerial vehicles, especially , lightweight, flapping-wing robots inspired by insects and birds, are being explored. Various studies on aerodynamics, kinematics, and control have reported interesting features, such as high maneuverability and efficiency with flapping wings [1, 2, 3, 4]. This study focuses on a hover-capable, tailless flapping-wing robot.
Various insect-inspired robots equipped with electromagnetic motors have been proposed [5, 6, 7, 8, 9]. In particular, a 28.2-g untethered battery-powered robot that can move freely in all directions and hover was demonstrated for the first time [8]. However, robots driven by electromagnetic motors require gears and transmission mechanisms to convert the rotating motion of the motor into high-speed reciprocating motion, and they are heavy, weighing 10 g or more. Extremely insect-scale robots weighing less than 1 g have been proposed [10, 11, 12, 13]. Their translational wing motion is generated directly by piezoelectric or dielectric elastomer actuators; gears are not required, but transmission may be necessary. However, in flight experiments, they are not battery-powered but are externally driven by wires (tethered flight). Therefore, controlled flights of battery-powered insect-scale robots weighing less than 10 g are still a challenge.
@@ I INTRODUCTION
Figure 1: Newly developed robot with four paired wings, eight piezoelectric actuators, a glass-fiber-reinforced plastic body, carbon-fiber-reinforced plastic legs, a wingspan of $110$ mm, and a weight of $1.52$ g.
Yaw control keeps on-board sensors pointing in the required direction and simplifies the horizontal controller design. However, yaw control using flapping wings is fundamentally more difficult than using rotating wings. Yaw torque generation using flapping wings is complex, as shown by many yaw control studies [14, 15, 16, 17, 18]. As demonstrated in [17], the speed ratio of the upstroke and downstroke of the piezoelectric actuator can be changed by adding a second-harmonic frequency to the fundamental frequency, generating yaw torque. This has been confirmed in experiments under conditions where the aircraft is fixed in the vertical direction. In [18], a yaw-controlled flight was demonstrated by a flapping-wing robot with two piezoelectric actuators, two wings, and a transmission. Here, the effect of phase shift in the second harmonic on the yaw torque was investigated and found to highly depend on manufacturing errors and other parameter changes. Therefore, the actuator drive system of the robot wing should be simple for yaw control. Furthermore, as the control signals for flapping wings become more complex, the circuits that generate the signals become more complex. For untethered, battery-powered flights, the weight and power consumption of the circuit must be minimized, and the robot should be able to operate with simple control signals.
@@
We previously investigated a direct-driven flapping wing using a piezoelectric unimorph actuator [19, 20]. It is simple because it has no displacement-enhancing structure (transmission-free structure), and the wings are driven by simply changing the voltage amplitude at the resonance frequency. Furthermore, to suppress the coupling effect between the wings and the body [12], we fabricated a robot with three paired-wing actuators [21]. However, it was unable to generate sufficient yaw torque, thus the yaw angle drifted. Therefore, we propose an insect-scale flapping-wing robot weighing 1.52 g (Fig. 1), which generates yaw torque using a tilting arrangement of four paired-wing actuators. The simpler structure of the system results in er manufacturing errors; however, it still has an offset in the lift force. In addition, it is difficult to accurately measure the instantaneous lift forces generated during the flapping motion. Therefore, we modeled the offset force and torque acting on the body and designed an adaptive control system. The major contributions of this study are summarized below:
@@
  • An adaptive controller is designed to account for the offset in the lift force.
  • We compare the performance of our controller with that of a linear quadratic integral (LQI) controller using numerical experiments and confirm the adaptability of our controller in the presence of an unknown lift offset.
@@
The remainder of this paper is organized as follows: In Section 2, the control-oriented model of the flapping-wing robot with four paired-wing actuators and the uncertainties of the lift force and torque are derived. An adaptive controller is proposed, and voltage amplitudes are calculated in Section 3. The performance of the LQI controller and our controller is compared for different lift force offsets through numerical experiments in Section 4. Then, an experimental evaluation of the proposed method is performed in Section 5. Finally, Section 6 concludes the paper.
II MODELING
Figure 2: Wing arrangement and direction of the lift forces
II-A Tilted Wing Arrangement
Fig. 2 shows the wing arrangement of the newly developed four-paired-wing robot. The wing arrangement is different from that of the three-paired-wing robot developed in our previous study [21]. Furthermore, to generate yaw torque, the four paired wings are tilted slightly by an angle $\beta$, which is the rotation angle around the Y-axis of the wing coordinate system.
The body force $f\in\mathbb{R}^{3}$ and torque ${\tau}\in\mathbb{R}^{3}$ are given by
$\begin{bmatrix}f\\ \tau\end{bmatrix}=Mf_{w}$ (1)
@@ II-A Tilted Wing Arrangement
$e_{i}=\begin{bmatrix}\cos\gamma_{i}\sin\beta_{i}\\ \sin\gamma_{i}\sin\beta_{i}\\ \cos\beta_{i}\end{bmatrix},p_{i}=\begin{bmatrix}a_{i}\\ b_{i}\\ 0\end{bmatrix}+\begin{bmatrix}\cos\gamma_{i}\cos\beta_{i}\\ \sin\gamma_{i}\cos\beta_{i}\\ -\sin\beta_{i}\end{bmatrix},$
$a_{1}=a_{4}=a,\quad a_{2}=a_{3}=-a,$
$b_{1}=b_{2}=b,\quad b_{3}=b_{4}=-b,$
$\beta_{1}=\beta_{2}=\beta_{3}=\beta_{4}=-\beta,$
$\gamma_{1}=\gamma,\quad\gamma_{2}=\pi-\gamma,\quad\gamma_{3}=\pi+\gamma,\quad\gamma_{4}=2\pi-\gamma,$
$a$ and $b$ are the lengths of the body along the X- and Y-axes, respectively, $\gamma$ is the rotation angle around the Z-axis of the body coordinate system, and $l$ is the distance between the mounting position of each wing and lift-force center.
II-B Unknown Offsets
The uncertainties of the body force and torque in (1) are expressed as offsets. They are due to unmodeled aerodynamic effects, manufacturing errors, and variations over time in system properties. The actual body force $f_{body}(=[f_{body,x},f_{body,y},f_{body,z}]^{\top})$ and torque $\tau_{body}$ are given by
@@ II-B Unknown Offsets
$\begin{bmatrix}{f}_{o}\\ {\tau}_{o}\end{bmatrix}\approx-{M}\delta_{f_{w}}-\delta_{M}{f}_{w}.$ (3)
@@ II-C Flight Dynamics
where $v(=[v_{x},v_{y},v_{z}]^{\top})$ is the translational velocity of the body in the global coordinate system, $\omega(=[\omega_{B,x},\omega_{B,y},\omega_{B,z}]^{\top})$ is the angular velocity of the body in the body coordinate system, $m$ and ${J}\in\mathbb{R}^{3\times 3}$ are the mass and inertia of the body, respectively, ${R}$ is the rotation matrix, ${g}=[0,0,g]^{\top}$, $g$ is the gravitational acceleration.
@@
II-D Control-Oriented Model
@@
where $T(>0)$ is a time constant and $f_{d,z}\in\mathbb{R}$ and ${\tau}_{d}\in\mathbb{R}^{3}$ are the demanded force and torque, respectively. Note that the lift force oscillates depending on the reciprocating motion of the wing. In this study, the oscillation component is not considered.
III FLIGHT CONTROLLER
The offset force and torque in (3) are unknown parameters. They are uncertain and cannot be measured. Furthermore, a change in the relationship between $f_{w}$ and $V$ occurs due to damage during use. Therefore, in this study, an adaptive controller is adopted.
III-A Velocity Control
@@ III-B Attitude Control
${\omega}_{d}$ $=$ $-G{K}_{\eta}{e}_{\eta},$ (11)
$\dot{{\omega}}_{d}$$\approx$ $-G{K}_{\eta}\dot{{\eta}},$
$\ddot{{\omega}}_{d}$$\approx$ $-G{K}_{\eta}\ddot{{\eta}},$
@@ III-C Vertical Control
$\frac{f_{d,z}}{m}$ $=$ $-k_{z}s_{z}+\ddot{z}_{r}+T\dddot{z}_{r}+g+\frac{\hat{f}_{o,z}}{m},$ (15)
$\dot{\hat{f}}_{o,z}$ $=$ $-\frac{\gamma_{z}}{m}s_{z},$ (16)
@@
III-D Lift-Force Demand and Flapping-Amplitude Control
@@
${f}_{w,d}=\begin{bmatrix}{M}_{1}(3,:)\\ {M}_{2}\end{bmatrix}^{-1}\begin{bmatrix}{f}_{d,z}\\ \boldsymbol{\tau}_{d}\end{bmatrix}.$ (19)
The flapping-wing actuator used in this study is driven by simply changing the flapping amplitude [19]. The relationship between the wing force $f_{w}$ and the flapping amplitude $V$ is modeled around the voltage at which the robot supports its own weight as follows [21]:
@@
As a result, the required flapping amplitude $V_{d}\in\mathbb{R}^{4}$ is derived from (19) and (20) as $V_{d}=h^{-1}(f_{w,d})$. Then, $V_{d}$ is multiplied by a sinusoidal wave and then applied to the robot. Note that this can be performed using a digital circuit, such as a pulse-width modulator or similar modulator, which is much simpler than a circuit that generates complex analog waveforms [22].
IV SIMULATION
The effectiveness of the proposed controller was validated through numerical simulations under unknown offsets. The proposed controller was compared with a conventional LQI controller derived from the linearization of (4) and (5), including the first-order lag of the wing forces.
IV-A Simulation Setup
Tables III and III show the model parameters of the robot and the control parameters in some numerical simulations, respectively. Here, the model parameters are not correct. Notably, $J$ is the value of a three-paired-wing robot in [21]. The weight is set at 2.0 g considering the untethered flapping-wing robot with an integrated battery [22].
TABLE I: Model Parameters

### v2

@@ Abstract
Insect-scale micro-aerial vehicles, especially lightweight, flapping-wing robots, are becoming increasingly important for safe motion sensing in spatially constrained environments such as living spaces. However, yaw control using flapping wings is fundamentally more difficult than using rotating wings. In this study, an insect-scale, tailless robot with four paired tilted flapping wings (weighing 1.52 g) was fabricated to enable simultaneous control of four states, including yaw angle. The controllability Gramian was derived to quantify the controllability of the fabricated configuration and to evaluate the effects of the tilted-wing geometry on other control axes. This robot benefits from the simplicity of directly driven piezoelectric actuators without transmission, and lift control is achieved simply by changing the voltage amplitude. However, misalignment or modeling errors in lift force can cause offsets. Therefore, an adaptive controller was designed to compensate for such offsets. Numerical experiments confirm that the proposed controller outperforms a conventional linear quadratic integral controller under unknown offset conditions. Finally, in a tethered and controlled flight experiment, yaw drift was suppressed by combining the tilted-wing arrangement with the proposed controller.
keywords:
Flapping-wing robot , Yaw control , Adaptive control , Micro aerial vehicle , Controllability
journal: Control Engineering Practice
1 Introduction
Unmanned aerial vehicles (UAVs) are increasingly being used to monitor and inspect large environments such as farms, roads, buildings, and bridges. They are also expected to be employed in spatially constrained environments such as living spaces. In confined spaces, UAVs must possess the ability to hover, exhibit high maneuverability, and be lightweight to ensure safety in the event of collisions. Therefore, flapping-wing robots, inspired by insects and birds, are more suitable for constrained spaces than the rotating-wing robots commonly used in UAVs.
High maneuverability and efficiency of flapping wings have been reported in various studies on the aerodynamics, kinematics and control of flapping-wing robots (Dickinson et al., 1999; Orlowski and Girard, 2012; Helbling and Wood, 2018; Phan and Park, 2019). While the incorporation of a tail enables passive flight and extends flight duration (Rosen et al., 2016; de Croon et al., 2016), it also diminishes maneuverability and complicates hover control due to the tail’s damping effect and sensitivity to external disturbances such as wind. In this paper, we consider a tailless configuration to achieve full controllability in a compact design, focusing on how simple flapping mechanisms can maintain yaw stability.
Several approaches to flapping-wing actuation have been studied and summarized in (Helbling and Wood, 2018; Phan and Park, 2019). Electromagnetic motors are a promising approach, and many robots have been proposed, often employing complex gear or linkage mechanisms to achieve reciprocating motion (de Croon et al., 2009; Keennon et al., 2012; Nguyen et al., 2015; Karásek et al., 2018; Phan et al., 2019). They are heavy, weighing $10$ g or more. In contrast, Piezoelectric actuators, by enabling direct flapping motion without mechanical transmission, make it possible to design extremely insect-scale robots weighing less than $1$ g (Wood, 2008; Ma et al., 2013; Ozaki and Hamaguchi, 2018; Chen et al., 2019). These flapping-wing actuations enable lightweight and inherently safe designs, but there are still challenges in achieving sufficient yaw control capability due to the limited yaw torque and coupling among control axes.
@@ 1 Introduction
Figure 1: Flapping-wing robot fabricated with four paired wings tilted by $\beta=20$ degree and $\gamma=60$ degree, equipped with eight piezoelectric actuators, a glass-fiber-reinforced plastic body, carbon-fiber-reinforced plastic legs, a wingspan of $110$ mm, and a weight of $1.52$ g.
Yaw control, which keeps onboard sensors aligned in the required direction, simplifies the design of the horizontal controller. However, generating yaw torque using flapping wings is complex (Helbling et al., 2014; Teoh, 2015; Fuller, 2019). As demonstrated in (Steinmeyer et al., 2019), yaw torque can be generated by adjusting the speed ratio between the upstroke and downstroke of piezoelectric actuators. However, this process is sensitive to manufacturing errors and parameter variations (Chukewad and Fuller, 2021). Therefore, yaw control with flapping wings should be achieved through a simple structure and a simple control signal.
@@
We previously investigated a direct-driven flapping wing using a piezoelectric unimorph actuator (Ozaki and Hamaguchi, 2018a, b). It is simple because it has no displacement-enhancing structure (transmission-free structure), and the wings are driven by simply changing the voltage amplitude at the resonance frequency. Furthermore, to suppress the coupling effect between the wings and the body (Ozaki and Hamaguchi, 2018), we fabricated a robot with three paired-wing actuators (Jimbo et al., 2020). However, because it could only control three states simultaneously, it was unable to generate yaw torque. Therefore, we propose an insect-scale flapping-wing robot weighing 1.52 g (Fig. 1), which generates yaw torque using a tilting arrangement of four paired-wing actuators. The effects of the tilted wings on controls other than yaw were examined using the controllability Gramian to evaluate the characteristics of the fabricated configuration. Furthermore, despite the simple structure, manufacturing errors can occur, resulting in a lift-force offset. In addition, it is difficult to accurately measure the instantaneous lift forces generated during the flapping motion. Therefore, we modeled the offset force and torque acting on the body and designed an adaptive control system. The major contributions of this study are summarized below:
  • 1.
@@
  • 2.
    The controllability of the fabricated wing arrangement is analyzed using a controllability Gramian. The advantages of the tilting angle for yaw control, as well as the limitations for other controls, are clarified.
  • 3.
    An adaptive controller is designed to account for lift force offset. Through numerical experiments, we compared its performance with a linear quadratic integral (LQI) controller and confirmed its adaptability to an unknown lift offset.
  • 4.
@@
The remainder of this paper is organized as follows: In Section 2, a control-oriented model of the flapping-wing robot with four paired-wing actuators is derived, and the uncertainties of the lift force and torque are formulated. The controllability Gramian is also derived to provide a basis for analyzing the control characteristics of the fabricated configuration. Section 3 presents the design of an adaptive controller that compensates for unknown lift-force offsets. In Section 4, numerical simulations are conducted to evaluate the controllability characteristics and to compare the performance of the proposed controller with that of a conventional LQI controller. Section 5 demonstrates a controlled flight experiment of the flapping-wing robot. Finally, Section 6 concludes the paper.
2 Modeling
2.1 Tilted-wing arrangement and yaw torque generation
Figure 2: Wing arrangement and lift force direction
Fig. 2 shows the wing arrangement of the newly developed four-paired-wing robot. The wing arrangement is different from that of the three-paired-wing robot developed in our previous study (Jimbo et al., 2020). Furthermore, to generate yaw torque, the four paired wings are tilted slightly by an angle $\beta$, which is the rotation angle around the Y-axis of the wing coordinate system.
The body force $f(=[f_{x},f_{y},f_{z}]^{\top})$ and torque ${\tau}(=[\tau_{x},\tau_{y},\tau_{z}]^{\top})$ are given by
$\begin{bmatrix}f\\ \tau\end{bmatrix}=Mf_{w},$ (1)
@@ 2.1 Tilted-wing arrangement and yaw torque generation
$e_{i}=\begin{bmatrix}\cos\gamma_{i}\sin\beta_{i}\\ \sin\gamma_{i}\sin\beta_{i}\\ \cos\beta_{i}\end{bmatrix},p_{i}=\begin{bmatrix}a_{i}+l\cos\gamma_{i}\cos\beta_{i}\\ b_{i}-l\sin\gamma_{i}\cos\beta_{i}\\ -l\sin\beta_{i}\end{bmatrix}$,
are the lengths of the body along the X- and Y-axes, respectively, $\beta_{1}=\beta_{2}=\beta_{3}=\beta_{4}=-\beta$, $\gamma_{1}=\gamma,\gamma_{2}=\pi-\gamma,\gamma_{3}=\pi+\gamma,\gamma_{4}=2\pi-\gamma$, $\gamma$ is the rotation angle around the Z-axis of the body coordinate system, and $l$ is the distance between the mounting position of each wing and lift-force center. The generated yaw torque $\tau_{z}$ is expressed as
$\tau_{z}=\sum_{i=1}^{4}\left(a_{i}\sin\gamma_{i}-b_{i}\cos\gamma_{i}\right)f_{i}\sin\beta_{i}.$
Note that when $\beta_{i}=0(i=1,\cdots,4)$, $\tau_{z}=0$.
2.2 Representation of unknown offsets
The body force and torque in (1) are challenging to model accurately due to unmodeled aerodynamic effects, manufacturing errors, and time-dependent variations in system properties. Piezoelectric materials also exhibit characteristic variations, with lift force variations of approximately 10–20 $\%$ in the paired wings fabricated in (Jimbo et al., 2020). A simple drive structure can mitigate these uncertainties, but they cannot be completely ignored. Therefore, the uncertainties in body force and torque are expressed as offsets. The actual body force $f_{body}(=[f_{body,x},f_{body,y},f_{body,z}]^{\top})$ and torque $\tau_{body}$ are given by
@@ 2.2 Representation of unknown offsets
$\begin{bmatrix}{f}_{o}\\ {\tau}_{o}\end{bmatrix}\approx-\delta_{M}{f}_{w}-{M}\delta_{f_{w}}.$ (3)
@@ 2.3 Flight dynamics
where $v(=[v_{x},v_{y},v_{z}]^{\top})$ is the translational velocity of the body in the global coordinate system, $\omega(=[\omega_{B,x},\omega_{B,y},\omega_{B,z}]^{\top})$ is the angular velocity of the body in the body coordinate system, $m$ and ${J}\in\mathbb{R}^{3\times 3}$ are the mass and inertia of the body, respectively, R{R} is the rotation matrix, $0,0,g]^{\top}, gg$ is the gravitational acceleration.
@@
2.4 Control-oriented formulation
@@
where $T(>0)$ is a time constant and $f_{d,z}\in\mathbb{R}$ and ${\tau}_{d}\in\mathbb{R}^{3}$ are the demanded force and torque, respectively. Note that the lift force oscillates due to the reciprocating motion of the wing. In this study, we focus on the steady lift force, ignoring high-frequency oscillations. While this simplification may limit control accuracy and design flexibility, it streamlines design, implementation, and maintenance while avoiding unexpected resonance.
2.5 Controllability
From (5) and (8), we consider the continuous-time linear time-invariant system
$\dot{\xi}=A\xi+Bf_{w},$ (11)
where $\xi=[mv_{z},(J\omega)^{\top}]^{\top}\in\mathbb{R}^{4}$, $A={\mathbf{0}}\in\mathbb{R}^{4\times 4}$, and $B=[M_{1}(3,:)^{\top},M_{2}^{\top}]^{\top}\in\mathbb{R}^{4\times 4}$. This equation is weight-compensated and ignores the offset and centrifugal Coriolis forces.
For the driftless system of (11), the controllability Gramian at time $t_{\rm f}$ is expressed as
$W_{c}(t_{\rm f})=\int_{0}^{t_{\rm f}}{\rm e}^{At}BB^{\top}{\rm e}^{A^{\top}t}dt=t_{\rm f}BB^{\top}.$
If $W_{c}(t_{\rm f})$, i.e. $BB^{\top}$ is a nonsingular matrix, a state transition to an arbitrary target can be realized at time $t_{\rm f}$. Then, the minimum input energy is expressed in quadratic form using the target state and the inverse of $W_{c}(t_{\rm f})$ (Kailath, 1980). As $BB^{\top}$ is singular for $B\in\mathbb{R}^{4\times 3}$ for the robot with three paired wings in (Jimbo et al., 2020), controlling four states of $\xi$ simultaneously requires a robot with four paired wings.
3 Flight controller
The offset force and torque in (3) are unknown parameters. They are uncertain and cannot be measured. Furthermore, a change in the relationship between $f_{w}$ and $V$ occurs due to damage during use. Therefore, an adaptive controller is adopted in this study.
3.1 Velocity control
@@ 3.2 Attitude control
${\omega}_{d}=-G{K}_{\eta}{e}_{\eta},\ \ \ \dot{{\omega}}_{d}\approx-G{K}_{\eta}\dot{{\eta}},\ \ \ \ddot{{\omega}}_{d}\approx-G{K}_{\eta}\ddot{{\eta}}$ (12)
@@ 3.3 Vertical control
$f_{d,z}/m$ $=$ $-k_{z}s_{z}+\ddot{z}_{r}+T\dddot{z}_{r}+g+\hat{f}_{o,z}/m,$ (16)
$\dot{\hat{f}}_{o,z}$ $=$ $-\gamma_{z}s_{z}/m,$ (17)
@@
3.4 Lift-force demand and flapping-amplitude control
@@
${f}_{w,d}=\begin{bmatrix}{M}_{1}(3,:)\\ {M}_{2}\end{bmatrix}^{-1}\begin{bmatrix}{f}_{d,z}\\ \bm{\tau}_{d}\end{bmatrix}$. (20)
The flapping-wing actuator used in this study is driven by simply changing the flapping amplitude (Ozaki and Hamaguchi, 2018a). The relationship between the wing force $f_{w}$ and the flapping amplitude $V$ is modeled around the voltage at which the robot supports its own weight as follows (Jimbo et al., 2020):
@@
As a result, the required flapping amplitude $V_{d}\in\mathbb{R}^{4}$ is derived from (20) and (21) as $V_{d}=h^{-1}(f_{w,d})$. Then, $V_{d}$ is multiplied by a sinusoidal wave and then applied to the robot. Note that this can be performed using a digital circuit, such as a pulse-width modulator or similar modulator, which is much simpler than a circuit that generates complex analog waveforms (Ozaki et al., 2023).
4 Simulation
The effectiveness of the proposed controller was validated through numerical simulations under unknown offsets. Before conducting the performance evaluations, the controllability of the flapping-wing robot fabricated with the tilting angles shown in Fig. 1 was examined using the controllability analysis described in Section 2.5. The proposed controller was then compared with a conventional LQI controller derived from the linearization of (4) and (5), including the first-order lag of the wing forces.
4.1 Controllability analysis
(a) Determinant
(b) Eigenvalues when $(a,b,l)=(20,5,40)$ $m$.
Figure 3: Controllability Gramian $BB^{\top}$
Fig. 3(a) shows the determinant of $BB^{\top}$ for varying $\beta$, $\gamma$, and the ratio of body lengths, $a/b$, at $ab=1$. A larger value of $a/b$ corresponds to a larger determinant of $BB^{\top}$ and a more controllable body. In this study, we set $(a,b)=(20,5)$ mm because $b=5$ mm is the smallest value feasible for manufacturing.
A single paired wing generates approximately $0.5$ gf at $120$ V (Jimbo et al., 2020), giving a total of $2.0$ gf for four pairs. The determinant peaks at $(\beta,\gamma)=(38,66)$ degrees and $a/b=4$ (Fig. 3(a)), where the effective lift force becomes $\cos\beta\times 2.0=0.79\times 2.0=1.58$ gf, corresponding to a lift-to-weight ratio of $1.04$ for the $1.52$ g robot. This result indicates that determinant maximization enhances controllability but leaves little lift margin. In contrast, the design setting of $\beta=20$ deg (white circle in Fig. 3(a)) yields a lift of approximately $1.88$ gf, corresponding to a lift-to-weight ratio of $1.24$. This ratio lies within a reasonable range, considering that the previous three-pair-wing robot achieved about $1.3$, and thus the design can be regarded as appropriate in terms of lift margin.
Directions in the state space corresponding to large eigenvalues of $BB^{\top}$ can be controlled with minimal input energy. Fig. 3(b) shows the eigenvalues of $BB^{\top}$ in the roll, pitch, yaw, and vertical directions. It can be observed that the vertical direction is the most controllable, while the yaw direction is the least controllable. In addition, prioritizing the yaw direction reduces the controllability in the vertical and pitch directions. For the selected value of $\gamma=60$ deg (white circle in Fig. 3(b)), the roll and pitch directions are well balanced, although the controllability in the yaw direction remains relatively weak.
Table 1: Plant Parameters